Please help!
I need to find the solution and graph for:
5x-6y=-25
3x+28=y
Please show me how to do this.
Thank you

Answer by stanbon(75887)

5x-6y=-25
3x+28=y
---------------
Substitute for "y" and solve for "x":
5x -6(3x+28) = -25
5x -18x -168 = -25
-13x = 143
x = -11
=================
Cheers,
Stan H.